HUBBLE SPACE TELESCOPE DAILY REPORT #1974 PERIOD COVERED: 0000Z (UTC) 09/30/97 - 0000Z (UTC) 10/01/97 Daily Status Report as of 274/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7693 (NICMOS Pupil Transfer Function) The Near Infrared Camera (Camera Two) was used to make observations of P330E-Cal in order to perform a pupil transfer function observations. The output of this program will be a two-dimensional surface fit that will give the large scale corrections needed to flat-field the images to an accuracy of about 2%. Also, changes in the PSF will be measured as a function of position within the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.5 Completed STIS/CCD 7276 (Synoptic Monitoring of Seasonal Phenomena on Mars) The Space Telescope Imaging Spectrograph (CCD) was used to make seasonal ultraviolet imaging of the planet Mars. This season includes the formation of the north polar cap and hood system, a process which is not well documented; the start of the classic dust storm season on Mars, when large dust storms have often occurred; and an active cycle of small dust storms in Hellas. These observations will complete a period of continuous monitoring of Mars which commenced in Cycle 6. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.7 Completed NIC/1/2/3 7474 (Searching for Low Mass Stars: the Mass Function at the Hydrogen Burning Limit) The Near Infrared Camera (Cameras One, Two, and Three) was used to look at the unprecedented IR imaging capabilities of NICMOS to study the low-mass stars at the end of the main sequence in four nearby globular clusters spanning a wide range of metallicity, precisely at the wavelength corresponding to their maximum emission {~1.1 Mum}. These observations would complete our ongoing investigation by making it possible, for the first time, to reliably count stars right down to and beyond the main sequence H-burning limit, thus leaving no doubt as to whether the IMF indeed drops there or starts rising once again close to the limit. The observation completed as planned with no problems. 1.8 Completed Two Sets of STIS/CCD 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Four Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Three Sets of STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 1.11 Completed Two Sets of NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-6384. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 6134 (Investigating the Nature of Circumstellar Matter in Young Clusters) The WF/PC-2 was used to take broad band images of several young star clusters associated with bright reflection nebulae in an attempt to detect circumstellar (CS) material about the stars in these clusters. The HST data will be combined with ground based near-IR and millimeter data. The combined data is expected to determine the frequency distribution of CS material and disks in undisturbed young clusters, the size and mass distribution of CS nebulae, and the correlation of these properties with stellar luminosity. The observations were completed as scheduled, and no problems were reported. 1.13 Completed NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one-orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 9 As described in HSTAR 6179, the acquisition at 274/053347Z defaulted to fine lock on GFS-1 only. Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: A STIS MCE-2 reset occurred at 273/185915Z, outside the SAA area. MAMA2 high voltage was off at the time of the reset. STIS MAMA recovery (MCE-2) was accomplished via operations request at 273/2300Z. No STIS proposal was impacted. HSTAR 6177 documents the event. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH